# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v2

@@ Abstract
Differentially Private Stochastic Gradient Descent (DPSGD) is widely used to protect sensitive data during the training of machine learning models, but its privacy guarantees often come at the cost of model performance, largely due to the inherent challenge of accurately quantifying privacy loss. While recent efforts have strengthened privacy guarantees by focusing solely on the final output and bounded domain cases, they still impose restrictive assumptions, such as convexity and other parameter limitations, and often lack a thorough analysis of utility. In this paper, we provide rigorous privacy and utility characterization for DPSGD for smooth loss functions in both bounded and unbounded domains. We track the privacy loss over multiple iterations by exploiting the noisy smooth-reduction property and establish the utility analysis by leveraging the projection’s non-expansiveness and clipped SGD properties. In particular, we show that for DPSGD with a bounded domain, (i) the privacy loss can still converge without the convexity assumption, and (ii) a smaller bounded diameter can improve both privacy and utility simultaneously under certain conditions. Numerical results validate our results.
Machine Learning, ICML
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DPSGD) (Abadi et al., 2016) has emerged as the leading defense mechanism to protect personal sensitive data in training of machine learning models. However, achieving good performance with DPSGD often comes with a significant privacy cost. A fundamental question, therefore, is how to precisely quantify the privacy loss associated with DPSGD.
@@
Previous methods for quantifying privacy loss include strong composition (Dwork et al., 2010; Bassily et al., 2014; Kairouz et al., 2015), moments accountant (Abadi et al., 2016), Rényi Differential Privacy (RDP) (Mironov, 2017; Mironov et al., 2019), and Gaussian Differential Privacy (GDP) (Dong et al., 2022), along with several numerical composition methods (Koskela et al., 2020; Gopi et al., 2021). These methods primarily rely on composition theorems, assuming that all intermediate models are revealed during the training procedure, which leads to an overestimation of privacy loss. While numerical composition methods aim to tightly characterize the privacy loss, they still operate under this same assumption.
@@
To address this overestimation, recent works have focused solely on the privacy guarantees of the final output. For instance, the privacy amplification by iteration (Feldman et al., 2018) demonstrated that withholding intermediate results significantly enhances privacy guarantees for smooth and convex objectives. Building upon this, Chourasia et al. (2021) suggest that the privacy loss of DPGD, the full batch version of DPSGD, may converge exponentially fast for smooth and strongly convex objectives. Furthermore, results by Ye and Shokri (2022) as well as Ryffel et al. (2022) extended this analysis to assess the privacy loss of DPSGD, although both studies rely on the assumption of strong convexity.
@@
More recently, the work by Altschuler and Talwar (2022) and its extension (Altschuler et al., 2024) established a constant upper bound on privacy loss after a burn-in period for Lipschitz continuous and smooth convex losses over a bounded domain. However, this analytical result is limited by its reliance on the convexity assumption and strict restrictions on the Rényi parameter $\alpha$, which hinders its broader applicability. In order to relax several strong assumptions, Kong and Ribero (2024) provided an analysis of weakly-convex smooth losses in the case where data is traversed cyclically. Later, Chien and Li (2024) suggest precisely tracking the privacy leakage incurred before reaching the constant upper bound by solving an optimization problem. However, this result is formulated as an optimization problem rather than a closed-form expression, making it hard to operationalize. Notably, most recent methods necessitate double clipping of both gradients and parameters due to the additional bounded domain assumption. These methods, however, do not provide a thorough utility analysis or experimental results, leaving their practical performance and trade-offs underexplored.
@@
[…]
@@
Feldman et al. 2018 convex, $L$-smooth unbounded $\mathcal{O}\left(\frac{\alpha C^{2}}{b^{2}\sigma_{\text{DP}}^{2}}T\right)$ $\surd$
Altschuler & Talwar 2022 convex, $L$-smooth, $M$-Lipschitz bounded $\mathcal{O}\left(\frac{\alpha M^{2}}{n^{2}\sigma_{\text{DP}}^{2}}\min\left\{T,\frac{Dn}{\eta M}\right\}\right)$$\times$
$\eta\leq 2/L$, $b\leq n/5$, $\sigma_{\text{DP}}>8\sqrt{2}M/b$, $\alpha\leq\alpha^{*}(b/n,\frac{b\sigma_{\text{DP}}}{2\sqrt{2}M})$
Kong & Ribero 2024 $m$-weakly convex, $L$-smooth bounded $\mathcal{O}\left(\frac{\alpha}{\sigma_{\text{DP}}^{2}}(D\sqrt{1+2\eta m[1+\frac{m}{2(L+m)}]}+\frac{\eta C}{b})^{2}\right)$$\times$
$\eta\leq\frac{1}{2(m+L)}$
Chien & Li 2024 $L$-smooth or $(L,\lambda)$-Hölder continuous gradient bounded w/o analytical form $\times$
@@
Ours $L$-smooth bounded $\mathcal{O}\left(\frac{\alpha}{\sigma_{\textup{DP}}^{2}}(\frac{(1+\eta L)^{2}D^{2}}{\eta^{2}}+\frac{C^{2}}{nb})\right)$ $\surd$
@@ 1.2 Other Related Works
In 2014, Bassily et al. (2014) analyzed the optimal utility guarantees of DPSGD under the assumption of Lipschitz continuity, considering both convex and strongly convex cases. Then, based on the additional assumption of the gradient distribution, Chen et al. (2020) studied the convergence of DPSGD with gradient clipping (DPSGD-GC) and derived a utility bound for the non-convex setting. The work by Song et al. (2021) explored the convergence of DPSGD-GC for generalized linear models, noting that, in the worst case, the utility can remain constant relative to the original objective. Later, Fang et al. (2023) further refined this analysis for smooth and unconstrained problems, providing more precise convergence results. However, many of these studies fix a specific value for the clipping threshold $C$, which may be adjusted due to privacy requirements.
More recently, Koloskova et al. (2023) characterized the convergence guarantees for DPSGD-GC across various clipping thresholds $C$ in the non-convex setting. While this work provides valuable convergence insights for DPSGD-GC, recent privacy characterizations have introduced the need for double clipping—clipping both gradients and parameters—due to the additional assumption of bounded domains. The convergence analysis involving double clipping has not been thoroughly explored in the existing literature.
@@ 2 Preliminaries
In this section, we introduce the foundational concepts and definitions relevant to our analysis. We start with our notation, which will be used throughout this paper.
Notations. Let $\operatorname{Pr}[\cdot]$ denote the probability of a random event, and $\mathbb{P}_{\boldsymbol{\mu}}$ be the law of a random variable $\boldsymbol{\mu}$. We refer to two datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ as adjacent if they differ from each other only by adding or removing one data point.
@@ 2.1 Rényi Differential Privacy (RDP)
We first recall the formal definition of differential privacy (DP), which provides a standard framework to ensure that a model’s output remains almost unchanged when applied to two adjacent datasets that differ only in a single data entry.
@@
(Differential privacy (Dwork et al., 2006)). For $\epsilon\geq 0$, $\delta\in[0,1]$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is $(\epsilon,\delta)$-DP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, and for any subset of outputs $\mathcal{S}\subseteq\mathcal{Y}$, we have
@@
(Rényi divergence (Rényi, 1961)). For adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, a mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$, and an outcome $s\in\mathcal{Y}$, the Rényi divergence of a finite order $\alpha\neq 1$ between $\mathcal{M}(\mathcal{D})$ and $\mathcal{M}(\mathcal{D}^{\prime})$ is defined as
@@
(Rényi differential privacy (Mironov, 2017)). For $\alpha>1$, $\varepsilon\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ satisfies $(\alpha,\varepsilon)$-RDP if, for any pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, it holds that
@@
(From $(\alpha,\varepsilon)$-RDP to $(\epsilon,\delta)$-DP (Mironov, 2017)). If $\mathcal{M}$ is an $(\alpha,\varepsilon)$-RDP mechanism, it also satisfies $(\varepsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $10<\delta<$ 1.
Based on the assumption that intermediate training models are not revealed, privacy amplification by iteration (Feldman et al., 2018) substantially improved the privacy guarantee analysis. This analytical framework is grounded on the concept of shifted Rényi divergence, which is also useful for our analysis.
@@
(Shifted Rényi divergence (Feldman et al., 2018)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $z\geq 0$ and $\alpha>1$, the $z$-shifted Rényi divergence is defined as
@@
where $W_{\infty}(\cdot,\cdot)$ denotes the $\infty$-Wasserstein distance [^1] See Definition A.6. between two distributions.
The privacy amplification by iteration framework relies on two key lemmas, which we restate below, focusing specifically on Gaussian noise, which suffices for our purposes. First, Lemma 2.6 stated next will prove to be useful for analyzing shifted Rényi divergence between two distributions convolved with Gaussian noise.
@@
(Shift-reduction (Feldman et al., 2018)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two $d$-dimensional random variables. Then, for any $a\geq 0$ and $z\geq 0$, we have
@@
where $\boldsymbol{\zeta}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{d})$, and $\mathbb{P}_{\boldsymbol{\mu}}*\mathbb{P}_{\boldsymbol{\zeta}}$ denotes the distribution of the sum $\boldsymbol{\mu}+\boldsymbol{\zeta}$ with $\boldsymbol{\mu}$ and $\boldsymbol{\zeta}$ drawn independently.
Lemma 2.7.
(Contraction-reduction (Feldman et al., 2018)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Suppose $\phi$, $\phi^{\prime}$ are two contractive functions222 A function is said to be contractive if it is $1$-Lipschitz. and let $\sup_{\boldsymbol{x}}\|\phi(\boldsymbol{x})-\phi^{\prime}(\boldsymbol{x})\|\leq s$. Then, for any $z\geq 0$, we have
Altschuler et al. (Altschuler & Talwar, 2022; Altschuler et al., 2024) provide a mild generalization of this lemma, extending it to be suitable for random contraction maps.
2.2 DPSGD with Double Clipping
Here, we consider the DPSGD with both gradient clipping and parameter projection (Algorithm 1), termed as DPSGD-DC. This method begins with applying the SGD procedure using Gaussian perturbation for model updates. Then, the updated parameters are projected into a bounded domain $\mathcal{K}$ with diameter $D$. Specifically, the update procedure can be expressed as
with
where $\eta$ denotes the learning rate, $\boldsymbol{\zeta}_{t}\sim\mathcal{N}(\boldsymbol{0},\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ with noise scale $\sigma_{\text{DP}}$, and $\nabla\mathcal{L}_{B_{t}}\left(\boldsymbol{\theta}_{t};\mathcal{D}\right)$ is the clipped SGD gradient obtained from a mini-batch $\mathcal{B}_{t}$, i.e., $\nabla\mathcal{L}_{B_{t}}\left(\boldsymbol{\theta}_{t};\mathcal{D}\right)=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}(\nabla l_{\xi}(\boldsymbol{\theta}_{t}))$, with $b=\left|\mathcal{B}_{t}\right|$ and $\operatorname{clip}_{C}(\boldsymbol{x})=\boldsymbol{x}\cdot\min(1,\frac{C}{\|\boldsymbol{x}\|})$.
If $\mathcal{K}=\mathbb{R}^{d}$, it reduces to the vanilla DPSGD with gradient clipping, and we refer to it as DPSGD-GC.

### v3

@@ Abstract
Differentially Private Stochastic Gradient Descent (DPSGD) is widely used to protect sensitive data during the training of machine learning models, but its privacy guarantee often comes at a large cost of model performance due to the lack of tight theoretical bounds quantifying privacy loss. While recent efforts have achieved more accurate privacy guarantees, they still impose some assumptions prohibited from practical applications, such as convexity and complex parameter requirements, and rarely investigate in-depth the impact of privacy mechanisms on the model’s utility. In this paper, we provide a rigorous privacy characterization for DPSGD with general $L$-smooth and non-convex loss functions, revealing converged privacy loss with iteration in bounded-domain cases. Specifically, we track the privacy loss over multiple iterations, leveraging the noisy smooth-reduction property, and further establish comprehensive convergence analysis in different scenarios. In particular, we show that for DPSGD with a bounded domain, (i) the privacy loss can still converge without the convexity assumption, (ii) a smaller bounded diameter can improve both privacy and utility simultaneously under certain conditions, and (iii) the attainable big-$\mathcal{O}$ order of the privacy utility trade-off for DPSGD with gradient clipping (DPSGD-GC) and for DPSGD-GC with bounded domain (DPSGD-DC) and $\mu$-strongly convex population risk function, respectively. Experiments via membership inference attack (MIA) in a practical setting validate insights gained from the theoretical results.
Codehttps://github.com/HauLiang/DPSGD-DC
Extended versionhttps://arxiv.org/abs/2502.17772
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DPSGD) (abadi2016deep) has emerged as the leading defense mechanism to protect personal sensitive data in training of machine learning models. However, achieving good performance with DPSGD often comes with a significant privacy cost. A fundamental question, therefore, is how to precisely quantify the privacy loss associated with DPSGD.
@@
Previous methods for quantifying privacy loss include strong composition (dwork2010boosting; bassily2014private; kairouz2015composition), moments accountant (abadi2016deep), Rényi Differential Privacy (RDP) (mironov2017renyi; mironov2019r), and Gaussian Differential Privacy (GDP) (dong2022gaussian), along with several numerical composition methods (koskela2020computing; gopi2021numerical). These methods primarily rely on composition theorems, assuming that all intermediate models are revealed during the training procedure, which leads to an overestimation of privacy loss. While numerical composition methods aim to tightly characterize the privacy loss, they still operate under this same assumption.
@@
To address this overestimation, recent works have focused solely on the privacy guarantees of the final output. For instance, the privacy amplification by iteration (feldman2018privacy) demonstrated that withholding intermediate results significantly enhances privacy guarantees for smooth and convex objectives. Building upon this, chourasia2021differential suggest that the privacy loss of DPGD, the full batch version of DPSGD, may converge exponentially fast for smooth and strongly convex objectives. Furthermore, results by ye2022differentially as well as ryffel2022differential extended this analysis to assess the privacy loss of DPSGD, although both studies rely on the assumption of strong convexity.
@@
More recently, the work by altschuler2022privacy and its extension (altschuler2024privacy) established a constant upper bound on privacy loss after a burn-in period for Lipschitz continuous and smooth convex losses over a bounded domain. However, this analytical result is limited by its reliance on the convexity assumption and strict restrictions on the Rényi parameter $\alpha$, which hinders its broader applicability. In order to relax several strong assumptions, kong2024privacy provided an analysis of weakly-convex smooth losses in the case where data is traversed cyclically. Later, chien2024convergent suggest precisely tracking the privacy leakage incurred before reaching the constant upper bound by solving an optimization problem. However, this result is formulated as a complex optimization problem rather than a closed-form expression, making it hard to operationalize. Notably, most recent methods necessitate double clipping of both gradients and parameters due to the additional bounded domain assumption. These methods, however, do not provide a thorough utility analysis or experimental results, leaving their practical performance and trade-offs underexplored.
@@
Table 1: Comparison of the $(\alpha,\varepsilon)$-RDP guarantee and assumptions needed by different works for DPSGD, where $b$ is the batch size, $n$ is the dataset size, $\eta$ is the step size, $C$ is the gradient clipping norm bound, $D$ is the diameter of the parameter domain, $\sigma_{\text{DP}}$ is the noise scale, and $T$ is the number of iterations. $\ddagger$}” is exclusively suitable for cyclic data traversal cases. $\dagger$}” indicates that a tighter bound can be obtained under additional assumptions on the Rényi parameters. $\alpha^{*}(q,\sigma)$ is defined as the largest $\alpha$ that satisfies both $\alpha\leq K\sigma^{2}/2-2\log\sigma$ and $\alpha\leq\left(K^{2}\sigma^{2}/2-\log 5-2\log\sigma\right)/\left(K+\log(q\alpha)+1/(2\sigma^{2})\right)$, where $K=\log(1+1/(q(\alpha-1)))$.
@@
feldman2018privacy convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ unbounded $\mathcal{O}\left(\frac{\alpha M^{2}}{b^{2}\sigma_{\text{DP}}^{2}}T\right)$ $\surd$
altschuler2022privacy convex, $L$-smooth, $M$-Lipschitz, $\eta\leq 2/L$ bounded $\mathcal{O}\left(\frac{\alpha M^{2}}{n^{2}\sigma_{\text{DP}}^{2}}\min\left\{T,\frac{Dn}{\eta M}\right\}\right)$ $\times$
$b\leq n/5$, $\sigma_{\text{DP}}>8\sqrt{2}M/b$, $\alpha\leq\alpha^{*}(b/n,\frac{b\sigma_{\text{DP}}}{2\sqrt{2}M})$
kong2024privacy $m$-weakly convex, $L$-smooth, $\eta\leq\frac{1}{2(m+L)}$ bounded $\mathcal{O}\left(\frac{\alpha}{\eta^{2}\sigma_{\text{DP}}^{2}}(D\sqrt{1+2\eta m[1+\frac{m}{2(L+m)}]}+\frac{\eta C}{b})^{2}\right)$ $\times$
chien2024convergent $M$-Lipschitz unbounded w/o analytical form $\times$
$L$-smooth or $(L,\lambda)$-Hölder continuous gradient /bounded
@@
Ours $L$-smooth bounded $\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}\min\left\{T,\frac{(1+\eta L)^{2}nbD^{2}}{\eta^{2}C^{2}}\right\}\right)$ $\surd$
@@ 1.1 Contributions
To clearly illustrate the effectiveness of our proposed privacy bound, we here provide detailed comparisons with several prominent existing approaches, including: a) feldman2018privacy, (b) the combined analysis by mironov2017renyi and mironov2019r, c) altschuler2022privacy, and d) kong2024privacy. The detailed setting can be found in Appendix D.1. As shown in Figure 1, our analysis demonstrates strict improvement on all existing privacy bounds, except for altschuler2022privacy. The reason for which their bound appeared tighter is their additional assumptions—including convexity and more restrictive conditions (as summarized in Table 1)—while our analysis relies only on the smoothness of the loss function and far weaker assumptions.
Figure 1: Comparison of our theoretical $(\alpha,\varepsilon)$-RDP bound for DPSGD-DC with existing approaches. Detailed assumptions required by each method have been summarized in Table 1.
@@ 1.2 Other Related Works
In 2014, bassily2014private analyzed the optimal utility guarantees of DPSGD under the assumption of Lipschitz continuity, considering both convex and strongly convex cases. Then, based on the additional assumption of the gradient distribution, chen2020understanding studied the convergence of DPSGD with gradient clipping (DPSGD-GC) and derived a utility bound for the non-convex setting. The work by song2021evading explored the convergence of DPSGD-GC for generalized linear models, noting that, in the worst case, the utility can remain constant relative to the original objective. Later, fang2023improved further refined this analysis for smooth and unconstrained problems, providing more precise convergence results. However, many of these studies fix a specific value for the clipping threshold $C$, which may be adjusted due to privacy requirements.
More recently, koloskova2023revisiting characterized the convergence guarantees for DPSGD-GC across various clipping thresholds $C$ in the non-convex setting. While this work provides valuable convergence insights for DPSGD-GC, recent privacy characterizations have introduced the need for double clipping—clipping both gradients and parameters—due to the additional assumption of bounded domains. The convergence analysis involving double clipping has not been thoroughly explored in the existing literature.
@@ 2 Preliminaries
In this section, we introduce the foundational concepts and definitions relevant to our analysis. We start with our notations, which will be used throughout this paper.
Notations. Let $\operatorname{Pr}[\cdot]$ denote the probability of a random event, and $\mathbb{P}_{\boldsymbol{\mu}}$ be the law of a random variable $\boldsymbol{\mu}$. We refer to two datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ as adjacent if they differ from each other by adding or removing only one data point.
@@ 2.1 Rényi Differential Privacy (RDP)
We first recall the formal definition of differential privacy (DP) and RDP.
@@
(Differential privacy (dwork2006our)). For $\epsilon\geq 0$, $\delta\in[0,1]$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is $(\epsilon,\delta)$-DP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, and for any subset of outputs $\mathcal{S}\subseteq\mathcal{Y}$, we have
@@
(Rényi divergence (renyi1961measures)). For adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$, and an outcome $s\in\mathcal{Y}$, the Rényi divergence of a finite order $\alpha\neq 1$ between $\mathcal{M}(\mathcal{D})$ and $\mathcal{M}(\mathcal{D}^{\prime})$ is defined as
@@
(Rényi differential privacy (mironov2017renyi)). For $\alpha>1$, $\varepsilon\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ satisfies $(\alpha,\varepsilon)$-RDP if, for every pair of adjacent datasets, $\mathcal{D},\mathcal{D}^{\prime}\subseteq\mathcal{X}$, it holds that
@@
(From $(\alpha,\varepsilon)$-RDP to $(\epsilon,\delta)$-DP (mironov2017renyi)). If $\mathcal{M}$ is an $(\alpha,\varepsilon)$-RDP mechanism, it is also $(\varepsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $0<\delta<1$.
Based on the assumption that intermediate training models are not revealed, the technique privacy amplification by iteration substantially improves the privacy guarantee analysis (feldman2018privacy), which is grounded on the concept of shifted Rényi divergence, as follows.
@@
(Shifted Rényi divergence (feldman2018privacy)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $z\geq 0$ and $\alpha>1$, the $z$-shifted Rényi divergence is defined as
@@
where $W_{\infty}(\cdot,\cdot)$ denotes the $\infty$-Wasserstein distance. [^1] See Definition A.7 in Appendix A.
The privacy amplification by iteration utilizes the following lemma, which we frequently employ in the sequel.
@@
(Shift-reduction (feldman2018privacy)). Let $\boldsymbol{\mu}$, $\boldsymbol{\nu}$ be two random variables. Then, for any $a\geq 0$ and $z\geq 0$, we have
@@
where $\boldsymbol{\zeta}$ is a multi-variate Gaussian random variable with zero mean and covariance given by $\sigma^{2}\boldsymbol{I}_{d}$, denoted by $\boldsymbol{\zeta}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{d})$; and $\mathbb{P}_{\boldsymbol{\mu}}*\mathbb{P}_{\boldsymbol{\zeta}}$ denotes the distribution of the sum $\boldsymbol{\mu}+\boldsymbol{\zeta}$ with $\boldsymbol{\mu}$ and $\boldsymbol{\zeta}$ being drawn independently.
2.2 DPSGD with Double Clipping (DPSGD-DC)